Monday, January 29, 2018

Update.

Sorry for the lack of updates recently, we had an extended microlensing season this year, our last microlensing target for the year was in Taurus and we took observations of that target throughout November, December and early January, we encountered all the usual challenges with frosting on the chip causing some frustration and missed nights, also during November we had a broken computer network so observations had to be taken using the computer in the dome which is never ideal, the target in Taurus was also fairly low in the sky for us, so not sure if our observations will be useful or not, although Subo Dong told me he was giving our data to his student to work on so we can only hope he can extract some useful data out of it and that Kumeu Observatory will get a mention in any papers written, we did have a pretty good run of weather during November and December so we managed to monitor the target with a daily cadence on almost every available clear night, including a lovely clear Christmas morning.

January hasn't been as good weather wise so far, it's been very hot with temperatures in the 18-20 degree range at night, our SBIG CCD can't cool down to -20 in these conditions, also means that the mosquito population has exploded so we either sweat it out inside the observatory or get eaten alive outside, work on the repairs to the dome have commenced with Peter constructing some scaffolding on the outside of the observatory, we also had an issue with the shutter when the elastic guide rope snapped one night when Jonathan Green was opening up giving him a fright.


<- The Dome scaffolding in place ready for the repair work.


Now that the microlensing observing campaign has finished, Jonathan Green, Steve Hennerley and Andrew Buckingham headed out in the daytime to take off the "Nustrini" C14 Schmidt-Cassegrain telescope and replaced it with a Astronz GSO 12' Ritchey–Chr├ętien telescope, we are going to have to find an adapter for the Optec Temperature Compensating Focuser and we will have to spend a few clear nights working to get everything setup and operational again, if the tests from the new optical tube are positive we may ask the council to purchase either a 14' or a 16' Ritchey–Chr├ętien telescope for the observatory as the "Nustrini" is really getting quite old now and is due for retirement.



We also repaired the elastic guide rope for the shutter and replaced the last two old filters in the filter wheel of the SBIG CCD camera so we now have an all new LRGB set as well as a new Wratten 12 filter, with the councils consent Jonathan Green and Steve Hennerley also removed the bunk beds downstairs to make more room for an observation workstation, we are very happy to have done this as on removal we found Rat feces and urine marks in both beds and really those beds were a health hazard, there is also evidence that Rats have been getting into the roof of the observatory but with a pig farm on one side and a macadamia nut farm on the other we are probably always going to have a rodent problem, all we can do is set rat baits and hope to control the problem as best we can.


<- The downstairs room with bunk beds removed, plenty of space now for an observation workstation to be installed



So we still have plenty to do out at Kumeu, Andrew Buckingham also has a QHY ColdCMOS camera for us to try out although it's yet to arrive in New Zealand, we will need to work on the RC's collimation, do another pointing model, take flats, work out our pixel scale, field of view etc etc so plenty to keep us busy over the coming months, bring on the clear skies!

Posted by Jonathan Green

No comments:

Post a Comment

Feel free to add comments - all comments are moderated, so please allow some time for them to show up.

Abusive, commercial, or junk comments are never shown!

Time Zone

All times on this page are in New Zealand Daylight Time in winter GMT +12 or summer GMT +13.